The probability that a patient recovers from a rare blood disease is 0.4. If 15 randomly chosen people are known to have contracted this disease, what is the probability that



(a) exactly 8 survive?



(b) at most 2 survive?



(c) less than 2 survive?



(d) at least 13 survive?



(e) more than 13 survive?



(f) between 5 and 8, exclusive, survive?



(g) between 5 and 8, inclusive, survive?



2. From the problem in number 1, how many are expected to survive from the 15 patients?

1
Expert's answer
2021-12-09T03:20:10-0500

Let "X=" the number of patients who survived: "X\sim Bin (n, p)."

Given "n=15, p=0.4, q=1-p=1-0.4=0.6."

(a)


"P(X=8)=\dbinom{15}{8}(0.4)^8(0.6)^{15-8}"

"=0.11805577445376\approx0.118056"

(b)


"P(X\leq 2)=P(X=0)+P(X=1)+P(X=2)"

"=\dbinom{15}{0}(0.4)^0(0.6)^{15-0}+\dbinom{15}{1}(0.4)^1(0.6)^{15-1}"

"+\dbinom{15}{2}(0.4)^2(0.6)^{15-2}"

"=0.02711400077216\approx0.027114"



(c)


"P(X< 2)=P(X=0)+P(X=1)"

"=\dbinom{15}{0}(0.4)^0(0.6)^{15-0}+\dbinom{15}{1}(0.4)^1(0.6)^{15-1}"

"=0.005172034830336\approx0.005172"

(d)


"P(X\geq 13)=P(X=13)+P(X=14)+P(X=15)"

"=\dbinom{15}{13}(0.4)^{13}(0.6)^{15-13}+\dbinom{15}{14}(0.4)^{14}(0.6)^{15-14}"

"+\dbinom{15}{15}(0.4)^{15}(0.6)^{15-15}"

"=0.000278904438784\approx0.000279"



(e)


"P(X>13)=P(X=14)+P(X=15)"

"=\dbinom{15}{14}(0.4)^{14}(0.6)^{15-14}+\dbinom{15}{15}(0.4)^{15}(0.6)^{15-15}"

"=0.000025232932864\approx0.000025"

(f)


"P(5<X<8)=P(X=6)+P(X=7)"

"=\dbinom{15}{6}(0.4)^{6}(0.6)^{15-6}+\dbinom{15}{7}(0.4)^{7}(0.6)^{15-7}"

"=0.38368126697472\approx0.383681"



(g)


"P(5\leq X\leq 8)=P(X=5)+P(X=6)"

"+P(X=7)+P(X=8)="

"=\dbinom{15}{5}(0.4)^{5}(0.6)^{15-5}+\dbinom{15}{6}(0.4)^{6}(0.6)^{15-6}"

"=\dbinom{15}{7}(0.4)^{7}(0.6)^{15-7}+\dbinom{15}{8}(0.4)^{8}(0.6)^{15-8}"

"=0.687674886193152\approx0.687675"



2.


"E(X)=np=15(0.4)=6"